A company that manufactures computer chips wants to use a multiple regression model to study the effect that the variables xdaily production volume (in 1000s of units) x2 daily amount of time involved in production (in hours) have on total daily production cost (in 1000s of dollars). If a regression model is estimated using 141 observations on each of these variables, what are the dearees of freedom (df) for the regression sum of squares, error sum of squares, and the total sum of squares? df for the regression sum of squares df for the error sum of squares df for the total sum of squares
Answer #1

Number of predictors = k = 2

Total observations on each of these variables = 141

df for the regression sum of squares=k=2

df for the error sum of squares=140-2=138

df for the total sum of squares =141-1=140